OpenStudy (anonymous):

Assume that it costs approximately dollars to manufacture x Xbox 360s in a day. C(x) = 62,500 + 550x + 0.01x^2 How many Xboxes should be manufactured in order to minimize average cost?

OpenStudy (anonymous):

minimum at the first coordinate of the vertex, which is always \(-\frac{b}{2a}\) in your example it is \(-\frac{550}{2\times (0.01)}\) which unfortunately is negative is there a typo in the problem?

OpenStudy (anonymous):

I dont think so, that is how the problem is written. Is there another way I could try to solve this problem?